A cell of electric motive force 2 volts supplies a current of 0.8 ampere through a cell of resistance 2 ohms . calculate the internal resistance of the cell
1
Expert's answer
2020-10-05T10:58:05-0400

Solution

Given data are

EMF of cell(E) =2V

Resistance (R) =2"\Omega"

Current (I) =0.8 A

So potential difference

V= IR="0.8\times2=1.6V"


Internal resistance of cell is given by "r=\frac{E-V}{I}=\frac{2-1.6}{0.8}=0.5\Omega"

Therefore internal resistance of cell is 0.5"\Omega."